# Tuning of Josephson junctions - the effects of depinning physics

### v2

@@ Abstract
Building more powerful quantum computers requires manufacturing processes with tight tolerances. To improve the tolerances on Josephson junctions, techniques to fine tune their properties after fabrication have been developed. Understanding how tuning techniques may physically modify the tunnel barrier of a Josephson junction is important and will enable these techniques to be optimised. We develop a model of junction tuning based on depinning theory to interpret a phase diagram of tuning rate. We extract the dependence on temperature, time-varying voltages and oscillation frequency. Using depinning theory we are able to show both why time-varying annealing potentials result in controlled junction tuning and how such protocols can be optimised. We examine how tuning changes the electrical breakdown of barriers and discrepancies between modeled and measured higher energy levels of transmon qubits.
@@
Quantum computers with high qubit count, low error rates and error correcting protocols promise new computational capabilities. Processors comprised of superconducting qubits are a leading platform for utility scale quantum computation [1, 2, 3] but building these at scale and with high quality poses many challenges, several of which lie in the processor manufacture. One such component which poses manufacturing difficulties is the Josephson junction (JJ). These are most often based on stacks of Al/AlOx/Al and are typically made by shadow angle evaporation [4] where a bottom layer of thin-film aluminium is oxidised before being coated by a second layer of aluminium. Uncontrolled variation in as-fabricated JJs currently limit qubit frequency assignment accuracy to 1 % [5, 6, 7], insufficiently precise for high fidelity 2Q gates in some gate modalities [8, 9]. Not only are junctions challenging to make with tight tolerance, but their properties change over time after their manufacture in a process known as aging [10] which complicates processor design. In order to address this variation a family of techniques which allow post-manufacture parameter adjustment have been developed including modifying the junction by laser irradiation [9, 11], electron beam irradiation [12] and alternating bias assisted annealing (ABAA) [13].
In this article we introduce a refined bias protocol using a lock-in amplifier, which naturally incorporates alternating bias [13]. The lock-in allows for fast measurements of resistance at variable frequencies allowing precise resistance targetting even with large tuning speed. We explore the voltage-temperature-frequency space of the tuning process and interpret the resultant phase diagram with a depinning model. We find good agreement between this model and the three-dimensional phase diagram. We compare tuned and untuned junctions measuring tunneling IV curves, breakdown voltages and Josephson harmonics [14].
@@
We build a setup shown schematically in Fig. 1 (a) where we connect a JJ in series with a load resistor, creating a potential divider, and connect this circuit between the output ports of a Zurich Instruments MFLI lock-in amplifier. Measuring the voltage drop across either of these elements allows us to infer the resistance of the JJ, the property we optimise in this process. We can control both the amplitude and the frequency of the AC voltage output from the lock-in amplifier. The sample containing JJs is placed on a Peltier heating stage to control the temperature of the JJ during this process.
@@
We measure the resistance of the JJ as a function of the lock-in frequency (Fig. 1 b) and find that we get constant values up to a frequency of 8 kHz, when the measured resistance increases. This increase in resistance is a measurement artefact due to RC filtering based on a computed cut-off frequency of 40 kHz, with most capacitance arising from coaxial cabling 111Using a series resistance of 15 k and a typical capacitance for BNC cables of 80pF/m. Considering 3m of cables, we compute an RC frequency of 44 kHz.. We confine our frequency operating window to frequencies of 1 kHz and below to ensure we infer the correct resistances of the JJs.
@@
where is the resistance at the start of the protocol and both and are fit parameters, with relating to the speed of tuning. We find this model fits all measured tuning curves well, and that tuning increases in speed with decreasing frequency. On some occasions, not shown in Fig. 1, the junction fails during the tuning process, becoming a short circuit. The rate of junction failures for a given process is another parameter we explore.
We show that the behaviour (speed and failure rate) of junctions undergoing voltage tuning can be understood qualitatively in terms of depinning theory. Depinning theory applies to a wide range of systems in which there is an interplay between disorder, temperature and applied stresses. This is seen in domain walls in magnets, the movement of vortices in superconductors, charge movement in Josephson junction arrays, dislocation movement, crack formation and even the formation of avalanches in sand piles [16, 17, 18, 19, 20].
The classic picture of depinning theory as applied to disordered media invokes two phases, a pinned static phase where the system is in a relaxed local minima and a sliding phase where local regions or domains within the system have a nonzero average velocity [21, 22]. In the pinned phase, any static or oscillating field can induce local microscopic changes, which can act as a dissipation channel [23], but not large scale movement or restructure. Above some threshold applied field (the ‘depinning threshold’) the movement of these local domains has some finite average velocity and can ultimately lead to avalanche behaviour.
We consider the junction dielectric to host domains which are trapped local minima. Applying sufficient field or temperature can move these domains, surmounting some effective potential barrier. Following Nattermann et al. [22], we define the depinning temperature where the thermal energy is equivalent to the barrier height. The depinning voltage is the point at which there is sufficient applied field for domains to continuously move through the material, resulting in an avalanche.
@@
Interpreting the junction tuning results in terms of depinning suggests a) if the creep region is larger for lower frequencies, and b) in the creep regime, there is a linear movement of defects without leading to full avalanche. This explains why lower frequencies result in greater resistance shifts for fixed values of temperature and voltage amplitude. At some low-frequency-limit we anticipate this trend stops, as domain movement saturates before the field polarity is reversed presenting a new energy landscape to the mobile domains. We can also understand the junction breakdown in terms of the depinning avalanche process for , which is both temperature and local microstate dependent.
@@
We continue to map out the phase diagram by measuring the tuning ‘speed’ parameter, in Eq. 1. Operating at 1 kHz, we vary the temperature and voltage amplitude of the tuning process. In Fig. 2 (c) we plot , finding a similar boundary, now at a lower voltage, consistent with the expected behaviour of . This process is then repeated at a different tuning frequency (103 Hz) with data shown in Fig. 2 (d). Again showing the expected increase in the creep regime with reducing frequency.
@@
Using the approach [22, 24] given in Eq. 2, along with a finite value of , we compare to the linear boundaries shown in Fig. 2c) and d). We find that while the direction of the experimental shift with bias frequency is expected, its magnitude is actually much smaller than depinning theory would suggest (or corresponds to unrealistically small – see supplementaary material). This suggests that the dynamics in the creep regime are more complicated than in previous studies of depinning.
The phase diagram in Fig. 2 (a) indicates sharp lines demarking regions of the phase diagram. However, increases smoothly with increasing voltage, starting at 0. Likewise, the onset of junction failure starts with a small subset of junctions failing, increasing to all junctions failing at higher voltage. Operationally these boundaries are ‘fuzzy’ which is consistent with the depinning process being dependent on the specific microstate of the amorphous oxide. As a result, considerations weighing process time against yield can be used to select a working point - even when the details of the underlying mechanism are not well understood.
@@
While the depinning picture provides a qualitative description of the observed dynamics, a quantitative model is needed to understanding the underlying microscopic processes. Time-dependent dielectric breakdown has been extensively studied in silica-based dielectrics [34, 35, 36], in magnetic tunnel junctions [37] and low-k dielectrics [38, 39], as well as AlOx films [40, 41, 42, 43]. In these materials, a variety of breakdown mechanisms have been studied, including charge trapping within the barrier, time modulation of this charge and electric field-induced distortion of bonds in the barrier. In particular studies of copper ion migration [38, 39] have considered the influence of both convective and diffusive flux of ions. These correspond to diffusion driven by the electric field itself, or by the stoichiometry gradient respectively. This suggests a possible two-process mechanism for breakdown and tuning of Josephson junctions. Namely that breakdown is enabled by the convective movement of atoms in the direction of the applied field, while the tuning of the resistance at lower fields is related to the redistribution of oxygen or defects in the barrier via lateral diffusion - mediated by the oscillating field. To test this possibility, we investigate the average barrier characteristics before and after tuning.
Following our recent work [44] we measure the electrical breakdown of junctions by ramping a DC voltage across these junctions from 0 to 1.7 V recording the current that flows through each junction as a function of voltage (Fig. 3 a). We fit the linear portion of this IV to extract the resistance of the JJ and fit the full IV to the Simmons model [45] fixing the area to that found by AFM measurements 8.86 nm2. At some voltage the junction fails and becomes short circuit which is recorded as the breakdown voltage (black star in Fig. 3 a). We compare two sets of junctions, distinguished by their tuning, with resistances shown in Fig. 3 (b). The first set of 20 junctions is tuned at 85 ∘C with a voltage of 0.92 V targeting 30 % total tuning and find an average tuning of 34.6 0.8 %. This systematic overshoot occurs due to continued resistance changes over time, after we have hit our target value, as shown in Ref. 33. During this process two junctions failed. The second set of 25 junctions are fabricated on the same chip, undergo the same 85 ∘C heating as the tuned junctions, but are not tuned by an alternating voltage. Despite undergoing no electrical tuning they are 4 % higher in resistance than the junctions before tuning due to the thermal processing they undergo, indicative of why a room temperature process is advantageous.
@@
where is the local thickness of the barrier and is the local dielectric strength of the barrier. Both of these values may, in principle, vary across the barrier [46, 47] and be modified by tuning. In the depinning model we describe above domains are mobile under the application of electric fields. These fields will be the strongest at the thinnest regions of the barrier which means that the the creep rate will also be greatest at these thinnest points. Such weak points in the barrier have been suggested to contribute to non-idealities [47, 14, 48] and breakdown [42] of the JJs. An approach that reduces the non-idealities associated to pin holes or weak points may have advantages in homogenising the current phase relations of junctions.
The higher-energy states of transmon qubits have recently been suggested to be sensitive to the nanostructure of the barrier in JJs [14]. In Fig. 3 (f) we show the results of measuring two coaxmon qubits [49] where one qubit has had a junction tuned by 34 % and the other has experienced the same thermal treatment but not been tuned. The results of Ref. 14 would suggest that by substantially modifying the barrier we might expect to see some difference to the frequencies of these higher levels. We follow the protocol from Ref. 50, treating the coaxmon as a qudit and tuning up a series of gates across the 6 energy levels shown schematically in Fig. 3 (f). We compute the Josephson energy and charging energy using qutip [51] by considering the first two transition frequencies and constraining them to match the results from the transmon Hamiltonian
where and are the charge and phase operators respectively and and are the charging and Josephson energies [52]. The tuned (untuned) coaxmon have junction resistances of 5802 (6021 ), charging energies of 166 MHz (168 MHz) and Josephson energies of 23.2 GHz (22.2 GHz). In Fig. 3 (f) we show the difference between the measured transition frequencies and the computed transition frequencies for the different qudit states for both a tuned and untuned coaxmon. We find there is no substantial difference between the two coaxmons.
@@
We show that tuning modifies, not only the resistance of the junctions, but also the barrier height in the junction and the breakdown voltage of the junctions. We contrast this to a measurement of ‘Josephson Harmonics’ which appear unchanged comparing tuned and untuned junctions. The interpretation in Ref. 14 suggests that deviations in these harmonics can be ascribed to the thinnest points of junctions. Taken together our measurements suggest that the depinning process modifying the junction resistance is not simply ‘filling in’ the thinnest points of barriers, but fundamentally modifies the junction microstructure.
@@ Acknowledgements.
- Van Damme et al. [2024] J. Van Damme, S. Massar, R. Acharya, T. Ivanov, D. Perez Lozano, Y. Canvel, M. Demarets, D. Vangoidsenhoven, Y. Hermans, J. Lai, et al., Advanced cmos manufacturing of superconducting qubits on 300 mm wafers, Nature , 1 (2024).
@@
- Note [1] Using a series resistance of 15 k and a typical capacitance for BNC cables of 80pF/m. Considering 3m of cables, we compute an RC frequency of 44 kHz.
@@
- Kennedy et al. [2025] O. W. Kennedy, K. G. Crawford, K. Shahbazi, and C. D. Shelly, Analysis of josephson junction barrier variation – a combined stem, breakdown and monte-carlo approach, arXiv preprint arXiv:2502.13072 (2025).

### v3

# A Depinning Model for Josephson Junction Tuning

@@ Abstract
Superconducting quantum computers leverage superconducting qubits, systems with well-defined energy levels, the spacing of which is determined by the properties of embedded Josephson junctions. Indeterminacy of Josephson junction manufacturing limits large scale superconducting circuits, however junction-by-junction tuning affords the possibility of an order of magnitude improvement of determinacy - although the origin of this tuning effect is uncertain. Understanding how tuning techniques may physically modify the tunnel barrier of a Josephson junction is important and will enable these techniques to be optimised. We develop a model of junction tuning based on depinning theory to interpret a phase diagram of tuning rate. We extract the dependence on temperature, time-varying voltages and oscillation frequency. Comparing the effect of junction tuning on room-temperature IVs, electrical breakdown and frequency of higher energy levels of transmon qubits indicates that the tuning process leaves the barrier thickness unchanged but modifies its composition.
@@
Quantum computers with high qubit count, low error rates and error correcting protocols promise new computational capabilities. Processors comprised of superconducting qubits are a leading platform for utility scale quantum computation [1, 2, 3] but building these at scale and with high quality poses many challenges. One component which poses manufacturing difficulties is the Josephson junction (JJ). These are most often based on stacks of Al/AlOx/Al and are typically made by shadow angle evaporation [4] where a bottom layer of thin-film aluminium is oxidised before being coated by a second layer of aluminium. State-of-the-art variation in as-fabricated JJs currently limits junction resistance accuracy to 2 % [5, 6, 7, 8, 9, 10]. The problem of junction variation will likely be exacerbated in larger wafers necessitated by larger processors and by the use of long-range couplers [11]. Simulations show that current manufacturing accuracy is insufficient to build large, high quality, quantum processors [10, 8, 5]. Improvements to frequency assignment enabled by tuning protocols, have been shown to improve processor yield experimentally [9]. Not only are junctions challenging to make with tight tolerance, but their properties change over time after their manufacture in a process known as aging [12] which complicates processor design. In order to address this variation a family of techniques which allow post-manufacture parameter adjustment have been developed including modifying the junction by laser irradiation [5, 13], electron beam irradiation [14] and alternating bias assisted annealing (ABAA) [15]. These techniques, referred to here as tuning, cause the resistance of junctions to increase only.
In this letter, we explore the voltage-temperature-frequency space, facilitated by a simplified lock-in tuning technique adapted from [15]. We interpret the resultant phase diagram using depinning theory which applies to a wide range of systems in which there is an interplay between disorder, temperature and applied stresses. This is seen in domain walls in magnets, the movement of vortices in superconductors, charge movement in Josephson junction arrays, dislocation movement, crack formation and even the formation of avalanches in sand piles [16, 17, 18, 19, 20]. We find good agreement between our depinning model and the three-dimensional phase diagram. We compare tuned and untuned junctions measuring tunneling IV curves, breakdown voltages and Josephson harmonics [21].
@@
We build a setup shown schematically in Fig. 1 (a) where we connect a JJ in series with a load resistor, creating a potential divider, and connect this circuit between the output ports of a Zurich Instruments MFLI lock-in amplifier. Measuring the voltage drop across either of these elements allows us to infer the resistance of the JJ, the property we optimise in this process. We can control both the amplitude and the frequency of the AC voltage output from the lock-in amplifier. The sample containing JJs is placed on a Peltier heating stage to control the temperature of the JJ during this process. JJs are fabricated in Ref [22], Dolan bridge junctions with an area of 84,000 nm2 are fabricated by shadow evaporations before and after a static oxidation step. An atomic force micrograph of a typical junction is shown in the supplemental materials [23] which includes Ref. [24].
@@
We measure the resistance of the JJ as a function of the lock-in frequency (Fig. 1 b) and find that we measure a constant value up to a frequency of 8 kHz, when the measured resistance increases. This increase in resistance is a measurement artefact due to RC filtering based on a computed cut-off frequency of 40 kHz, with most capacitance arising from coaxial cabling - a series resistance of 15 k (10 k load and 5 k junction resistance) together with 3 m of BNC cables (typical capacitance of 80 pF/m) gives an RC frequency of 44 kHz. We confine our frequency operating window to frequencies of 1 kHz and below to ensure we infer the correct resistances of the JJs.
@@
where is the resistance at the start of the protocol and both and are fit parameters, with relating to the speed of tuning. We find this model fits all measured tuning curves well, and that tuning increases in speed with decreasing frequency. On some occasions, not shown in Fig. 1, the junction fails during the tuning process, becoming a short circuit. The rate of junction failures for a given process is another parameter we explore. We show that the behaviour (speed and failure rate) of junctions undergoing voltage tuning can be understood qualitatively in terms of depinning theory.
The classic picture of depinning theory as applied to disordered media invokes two phases, a pinned static phase where the system is in a relaxed local minimum and a sliding phase where local regions or domains within the system have a nonzero average velocity [25, 26]. The domains of depinning theory are a general concept, which may sometimes be instantiated by magnetic domains specifically, but this commonality in terminology is not fundamental. For example, domains may similarly be regions of sand grains, or any other type of effective degree of freedom that ‘depins’ in a disordered system. In the pinned phase, any static or oscillating field can induce local microscopic changes, which can act as a dissipation channel [27], but not large scale movement or restructure. Above some threshold applied field (the ‘depinning threshold’) the movement of these local domains has some finite average velocity and can ultimately lead to avalanche behaviour.
We consider the junction dielectric to host domains which are trapped local minima. These domains may be 0D vacancies/interstitials, 1D chains of such vacancies, or 2D boundary regions of lower density amorphous material or something else - the depinning treatment here is agnostic to the underlying microphysics. Applying sufficient field or temperature can move these domains, surmounting some effective potential barrier. Following Nattermann et al. [26], we define the depinning temperature where the thermal energy is equivalent to the barrier height. The depinning voltage is the point at which there is sufficient applied field for domains to continuously move through the material, resulting in an avalanche.
@@
Interpreting the junction tuning results in terms of depinning suggests a) if the creep region is larger for lower frequencies, and b) in the creep regime, there is a linear movement of defects without leading to full avalanche. This explains why lower frequencies result in greater resistance shifts for fixed values of temperature and voltage amplitude. At some low-frequency-limit we anticipate this trend stops, as domain movement saturates before the field polarity is reversed, presenting a new energy landscape to the mobile domains. We can also understand the junction breakdown in terms of the depinning avalanche process for , which is both temperature and local microstate dependent.
@@
We continue to map out the phase diagram by measuring the tuning ‘speed’ parameter, in Eq. 1. Operating at 1 kHz, we vary the temperature and voltage amplitude of the tuning process. In Fig. 2 (c) we plot , finding a similar boundary, now at a lower voltage, consistent with the expected behaviour of . This process is then repeated at a different tuning frequency (103 Hz) with data shown in Fig. 2 (d), showing the expected increase in the creep regime with reducing frequency.
@@
Using the approach [26, 28] given in Eq. 2, along with a finite value of , we compare to the stable/creep boundaries shown in Fig. 2c) and d). We find that while the direction of the experimental shift with bias frequency is expected, its magnitude is actually much smaller than depinning theory would suggest (or corresponds to unrealistically small – see supplementary material [23]). This suggests that the dynamics in the creep regime are more complicated than in previous studies of depinning.
Fig. 2 (a) indicates sharp lines demarking regions of the phase diagram, however, at this stage it is not clear that these boundaries are true phase transitions. The speed parameter increases smoothly with increasing tuning voltage (the data is shown with alternate visualisation in the supplementary materials [23]). This may be indicative of a second order phase transition between the stable and creep regimes.
@@
While the depinning picture provides a qualitative description of the observed dynamics, a quantitative model is needed to understanding the underlying microscopic processes. Time-dependent dielectric breakdown has been extensively studied in silica-based dielectrics [37, 38, 39], in magnetic tunnel junctions [40] and low-k dielectrics [41, 42], as well as AlOx films [43, 44, 45, 46]. In these materials, a variety of breakdown mechanisms have been studied, including charge trapping within the barrier, time modulation of this charge and electric field-induced distortion of bonds in the barrier. In particular studies of copper ion migration [41, 42] have considered the influence of both convective and diffusive flux of ions. These correspond to diffusion driven by the electric field itself, or by the stoichiometry gradient respectively. This suggests a possible two-process mechanism for breakdown and tuning of Josephson junctions. Namely that breakdown is enabled by the convective movement of atoms from within the barrier in the direction of the applied field, while the tuning of the resistance at lower fields is related to the redistribution of oxygen or defects from within the barrier via diffusion - mediated by the oscillating field. To test this possibility, we investigate the average barrier characteristics before and after tuning.
Following our recent work [47] we measure the electrical breakdown of junctions by ramping a DC voltage across these junctions from 0 to 1.7 V recording the current that flows through each junction as a function of voltage (Fig. 3 a). We fit the linear portion of this IV to extract the resistance of the JJ and fit the full IV to the Simmons model [48] fixing the area to that found by AFM measurements 8.86 nm2. At some voltage the junction fails and becomes short circuit which is recorded as the breakdown voltage (black star in Fig. 3 a). We compare two sets of junctions, distinguished by their tuning, with resistances shown in Fig. 3 (b). The first set of 20 junctions is tuned at 85 ∘C with a voltage of 0.92 V targeting 30 % total tuning and find an average tuning of 34.6 0.8 %. This systematic overshoot occurs due to continued resistance changes over time, after we have hit our target value, as shown in Ref. 6. During this process two junctions failed. The second set of 25 junctions are fabricated on the same chip, undergo the same 85 ∘C heating as the tuned junctions, but are not tuned by an alternating voltage. Despite undergoing no electrical tuning they are 4 % higher in resistance than the junctions before tuning due to the thermal processing they undergo. This shows why a room temperature process (as shown in the supplementary materials [23]) is advantageous as heating chips above storage temperatures causes uncontrolled changes in junction properties alongside the intentional tuning process.
@@
where is the local thickness of the barrier and is the local dielectric strength of the barrier. Both of these values may, in principle, vary across the barrier [49, 50] and be modified by tuning. In the depinning model we describe above domains are mobile under the application of electric fields. These fields will be the strongest at the thinnest regions of the barrier which means that the the creep rate will also be greatest at these thinnest points. Such weak points in the barrier have been suggested to contribute to non-idealities [50, 21, 51] and breakdown [45] of the JJs. An approach that reduces the non-idealities associated to pin holes or weak points may have advantages in homogenising the current-phase relations of junctions.
The higher-energy states of transmon qubits have recently been suggested to be sensitive to the nanostructure of the barrier in JJs [21]. In Fig. 3 (f) we show the results of measuring two coaxmon qubits [52] where one qubit has had a junction tuned by 34 % and the other has experienced the same thermal treatment but not been tuned. The results of Ref. 21 would suggest that by substantially modifying the barrier we might expect to see some difference to the frequencies of these higher levels. We follow the protocol from Ref. 53, treating the coaxmon as a qudit and tuning up a series of gates across the 6 energy levels shown schematically in Fig. 3 (f). We compute the Josephson energy and charging energy using qutip [54] by considering the first two transition frequencies and constraining them to match the results from the transmon Hamiltonian
where and are the charge and phase operators respectively and and are the charging and Josephson energies [55]. The tuned (untuned) coaxmon have junction resistances of 5802 (6021 ), charging energies of 166 MHz (168 MHz) and Josephson energies of 23.2 GHz (22.2 GHz). In Fig. 3 (f) we show the difference between the measured transition frequencies and the computed transition frequencies for the different qudit states for both a tuned and untuned coaxmon. We find the relative difference in model/experiment discrepancies for tuned and untuned junctions to be small relative to lab-to-lab variation or even qubit-to-qubit variation from a single institute presented in Ref. [21] and therefore conclude that the tuning process has not substantially modified the Josephson harmonics in this instance. Statistical studies comparing multiple tuned and untuned qubits may allow small effects to be resolved in future studies.
@@
We show that tuning modifies, not only the resistance of the junctions, but also the barrier height in the junction and the breakdown voltage of the junctions. We contrast this to a measurement of ‘Josephson Harmonics’ which appear unchanged comparing tuned and untuned junctions. The interpretation in Ref. 21 suggests that deviations in these harmonics can be ascribed to the thinnest points of junctions. Taken together our measurements suggest that the depinning process modifying the junction resistance is not simply ‘filling in’ the thinnest points of barriers, but fundamentally modifies the junction microstructure.
@@ Acknowledgements.
- Muthusubramanian et al. [2024] N. Muthusubramanian, M. Finkel, P. Duivestein, C. Zachariadis, S. L. van der Meer, H. M. Veen, M. W. Beekman, T. Stavenga, A. Bruno, and L. DiCarlo, Wafer-scale uniformity of dolan-bridge and bridgeless manhattan-style josephson junctions for superconducting quantum processors, Quantum Science and Technology 9, 025006 (2024).
@@
- Zhang et al. [2022] E. J. Zhang, S. Srinivasan, N. Sundaresan, D. F. Bogorin, Y. Martin, J. B. Hertzberg, J. Timmerwilke, E. J. Pritchett, J.-B. Yau, C. Wang, et al., High-performance superconducting quantum processors via laser annealing of transmon qubits, Science Advances 8, eabi6690 (2022).
@@
- Du et al. [2025] Z. Du, S. Kan, S. Stein, Z. Liang, A. Li, and Y. Mao, Hardware-aware compilation for chip-to-chip coupler-connected modular quantum systems, arXiv preprint arXiv:2505.09036 (2025).
@@
- [23] See Supplemental Material at URL-will-be-inserted-by-publisher.
- Dobrovinskaya et al. [2009] E. R. Dobrovinskaya, L. A. Lytvynov, and V. Pishchik, Properties of sapphire, Sapphire , 55 (2009).
@@
- Kennedy et al. [2025] O. W. Kennedy, K. G. Crawford, K. Shahbazi, and C. D. Shelly, Analysis of josephson junction barrier variation: A combined electron microscopy, breakdown, and monte carlo approach, Physical Review Materials 9, 084803 (2025).